OpenStudy (anonymous):

NEED HELP FOR EXAM Give the probability distribution for the indicated random variable. A fair coin is tossed three times, and we count the number of "runs" that occur. A "run" is defined as a sequence of one or more of the same letter (e.g., HTH is 3 runs, TTH is 2 runs, and TTT is 1 run).

OpenStudy (anonymous):

Underlying distribution is binomial p=1/2, q=1/2. However, runs require three in a row, so p = (1/2)^3 and q=1-p, Further, HHH or TTT counts, so must double p to p=(1/2)^2

OpenStudy (anonymous):

so i have a table like this

OpenStudy (anonymous):

how do you find whats missing?

OpenStudy (anonymous):

@douglaswinslowcooper